What is Sn for the arithmetic series with the d= -4,An=27 and n=9?
A)418 B)360 C)451 D)387

Answer by richwmiller(17219)

Find first term
An = a1 + (n - 1)d
27 = a1 + (8)-4
27=a1-32
a1=59
Find sum
S = ½(a1 + an)n
S = ½(59 + 27)9
S = ½(86)9
S = 43*9
S=387
D)387